# Complex-Valued Holographic Radiance Fields

### v1

@@ 3. Background and Problem Definition
Before explaining our approach, we first introduce the key concepts and terminologies in this paper to define the research problem we are trying to solve. In this section, we explain several primary components to describe the research problem. These components are: 3D Gaussians for scene representation, hologram planes for wave recording, and camera parameters for view-dependent rendering.
@@
Inline with the approach of (Kerbl et al., 2023), we learn geometry and radiance of scenes through the optimization of 3D Gaussians primitives. We consider a 3D region in Euclidean space, $\Omega\subset\mathbb{R}^{3}$. A scene inside $\Omega$ is described using a set of $N$ oriented 3D Gaussians, and each Gaussian is denoted by $\mathcal{G}_{n}$. Typically, each $\mathcal{G}_{n}$ is parameterized by $k=(\mathbf{c}_{n},\mathbf{x}_{n},\mathbf{R}_{n},\mathbf{S}_{n},\boldsymbol{\alpha}_{n})$, where $\mathbf{c}_{n}\in\mathbb{R}^{3}$ represents color, $\mathbf{x}_{n}\in\mathbb{R}^{3}$ represents the center point of a Gaussian, $\mathbf{R}_{n}\in\mathbb{R}^{4}$ describes the rotation of a Gaussian, $\mathbf{S}_{n}$ represents scale along principal axes defined by $\mathbf{R}_{n}$, and $\boldsymbol{\alpha}_{n}$ represents the opacity. The geometrical shape of a 3D Gaussian is defined as:
[…]
where the covariance, $\Sigma$, is decomposed into a rotation matrix, $\mathbf{R}$, and a scaling matrix, $\mathbf{S}$:
[…]
We project these 3D Gaussians onto 2D planes to render images at a camera plane, a process also known as the differential splatting. The process of differential splatting involves applying the transformation matrix, $\mathbf{W}$, and the Jacobian projection function, $\mathbf{J}$, mapping 3D points to 2D image plane. The covariance in camera-space, $\Sigma^{\prime}$, is defined as:
@@
For each pixel, the color and opacity contributed by each Gaussian are computed by evaluating Eq. (2) under its learned parameters. The blending of $N$ ordered Gaussians overlapping a given pixel is defined as follows:
@@
We introduce our hologram plane as $P\in\mathbb{C}^{n_{x}\times n_{y}}$, where $n_{x},n_{y}\in\mathbb{Z}$. The center of $P$ is co-located with a camera pointing at the center of $\Omega$ located at $d$ distance away. We always position $P$ based on the camera pose, location and angles, $r=(x_{c},y_{c},z_{c},\phi_{x},\phi_{y},\phi_{z})$, in the Euler space. Each complex pixel $P_{i}$ has a pixel pitch, $\Delta x$. Given input $q=(r,\Delta x,n_{x},n_{y})$, the goal is to obtain the hologram plane by propagating the complex field generated from the Gaussian primitives to the hologram plane $P(q)$:
(6) $P(q)=U(q)*h,$
where $U(q)$ denotes the complex field rendered from the Gaussian primitives given camera parameters $q$, and $h$ is the spatial-domain wave propagation kernel derived from the inverse Fourier transform of the frequency-domain propagation function $H$ (Matsushima and Shimobaba, 2009; Zheng et al., 2024). A common propagation method is the band-limited , where the transfer function $H_{z}(f_{x},f_{y})$ encodes propagation by a distance $z$ in the frequency domain:
@@
In this paper, we also utilize band-limited as the propagation method for complex-valued holographic radiance fields modeling.
@@ 4. Method
In this section, we describe our method using the terms defined in Sec. 3. While classical radiance is defined as a scalar energy quantity ($\mathrm{W/(sr\cdot m^{2})}$) that does not encode phase, our formulation instead models a full complex-valued field, incorporating both amplitude and phase, to capture phenomena such as interference and diffraction. We use the term complex-valued holographic radiance fields to maintain conceptual continuity with prior 3D scene representations such as and , while reinterpreting radiance in a wave-optical context. This formulation bridges graphics and optics by embedding coherent light modeling within the familiar structure of radiance fields.
@@ 4.1. Complex-Valued Holographic Radiance Fields
We provide an overview of our method in Fig. 3. Given a camera pose, $r$, and a target hologram plane, $P$, we partition $\Omega$ into multiple planes in depth. At each depth plane, we render our learnable complex-valued 3D Gaussians into corresponding 2D projections. During this projection process, we utilize learnable probabilities to our 3D Gaussians that helps us determine if a 3D Gaussian contributes to that selected depth plane. Note that learning these probabilities in terms of their contributions at each depth plane help to inherently capture depth characteristics of a given scene. Once we determine the appropriate 2D projections for each layer, we propagate the complex field from each layer towards our camera to construct our $P$.
@@
For every 3D Gaussian $\mathcal{G}_{n}$ with center $\mathbf{x}_{n}$, we project its 2D projections to the ”most appropriate” plane $\Pi_{l}$, based on learnable probabilities. Let $\{\Pi_{1},\Pi_{2},\dots,\Pi_{L}\}$ be a set of discretized parallel planes in the volume, $\Omega$, positioned periodically along the optical axis in front of the view-dependent camera. Each plane $\Pi_{l}$ is separated from its neighbors by a fixed depth interval of $\Delta z$. The planes span a volume of depth, the center of the volume is placed at a propagation distance, $d$, from the camera plane. In this arrangement, the entire depth is symmetrically distributed as in the common literature, ranging from: $Z_{1}=d-\frac{(L-1)}{2}\Delta z$ to $Z_{L}=d+\frac{(L-1)}{2}\Delta z.$ Each plane $\Pi_{l}$ is sampled on a 2D grid that matches the physical resolution of a .
@@
Consider a multi-plane case where Gaussians are projected onto $L$ numbers of planes $\{\Pi_{l}\}_{l=1}^{L}$, we compute each 2D splat via perspective projection of 3D Gaussiani following Eq. (4),
[…]
where $u=(n_{x},n_{y})$ and $\mu=\mathbf{J}\,\mathbf{W}\,\mathbf{x}$. To model complex-valued holographic radiance fields, we have to modify the original definition of the 3D Gaussian. Specifically, each Gaussian is now parameterized by
(9) $k=(\mathbf{c}_{n},\mathbf{x}_{n},\mathbf{R}_{n},\mathbf{S}_{n},\boldsymbol{\alpha}_{n},\boldsymbol{\varphi}_{n},\mathbf{p}_{n}),$
where we add two extra parameters: $\boldsymbol{\varphi}_{n}\in\mathbb{R}^{3}$ to represent their inherent phase across wavelengths, and $\mathbf{p}_{n}\in\mathbb{R}^{L}$ to represent the plane assignment probabilities, where each element lies in $[0,1]$. Under this formulation, $\mathbf{c}_{n}$ denotes the inherent wave amplitude, replacing its original interpretation as color. The complex field of each 2D projections $U_{n}$ then is given by:
@@
When $L=0$, plane assignment reduces to a single plane, the $\mathbf{p}_{n}$ trivially equals 1.0, and the complex-valued holographic radiance field will generate 2D holograms. For the multi-plane case, we convert the pre-activation $\mathbf{p}^{\prime}_{n}$ to probabilities via arg-max followed by one-hot encoding, resulting in a hard assignment vector $\mathbf{p}_{n}$:
(11) $\mathbf{p}_{n}$ $=\mathrm{OneHot}(\arg\max_{l}\,\mathbf{p}^{\prime}_{n}),$
[…]
where $\mathbf{p}_{n,l}$ is the assignment probability of the $n$-th Gaussian to the $l$-th depth plane. However, the argmax operation is non-differentiable, which impedes gradient-based training. To address this, we employ the that allows gradient propagation through the discrete assignment:
[…]
Backward: $\frac{\partial\mathcal{L}}{\partial\mathbf{p}^{\prime}_{n}}=\frac{\partial\mathcal{L}}{\partial\mathbf{p}_{n}}\cdot\text{softmax}(\mathbf{p}^{\prime}_{n}/\tau),$
where $\tau$ is the temperature controlling the sharpness of the softmax approximation, in our case, we set $\tau=0.001$, which makes the softmax function close to argmax operation.
@@
The final complex field at each plane $\Pi_{l}$ is then defined as:
(13) $U_{\Pi_{l}}=\sum_{n=1}^{N}\mathbf{p}_{n,l}U_{n}\boldsymbol{\alpha}_{n}\prod_{j=1}^{n-1}\bigl{(}1-\boldsymbol{\alpha}_{j}\bigr{)}\mathbf{p}_{j,l}.$
This formulation ensures each complex-valued 3D Gaussian contributes exclusively to its assigned plane, allowing the framework to learn optimal plane assignments of the Gaussians during training. Unlike the alpha blending used in intensity-based approaches, which indiscriminately mixes RGB colors across all elements, our method creates plane-specific visibility by blending complex values (containing both amplitude and phase information) only among Gaussians assigned to the same depth plane to each pixel. Once each layer $\Pi_{l}$ is populated by assigned 2D projections, we use the transfer function $H$ to propagate $\Pi_{l}$ towards the hologram plane $P$, a process we named Forward Recording in our framework. This wave-based propagation extends the traditional intensity neural rendering in Eq. (5). Specifically, for every $\Pi_{l}$, a $(n_{x}^{(l)}\times n_{y}^{(l)})$ grid of complex samples, we compute:
@@
where $Z_{l}$ is the distance between depth plane $\Pi_{l}$ and the hologram plane $P$. The forward propagated field $U_{\Pi_{l}\to P}$ is then added with the contributions from the other planes to record the final complex 3D hologram:
@@
Since each plane is a 2D raster, the computation cost scales linearly with the number of layers $L$ and is dominated by s of size $n_{x}^{(l)}\times n_{y}^{(l)}$. After obtaining $P$, we propagate $P$ back to each depth plane. The back-propagated complex field $U_{P\to\Pi_{l}}$ at depth plane $\Pi_{l}$ is computed as:
@@
To further improve the image quality of the defocus region, we utilize two complementary loss functions. First, the reconstruction loss $\mathcal{L}_{recon}$ introduced by Kavakli et al. (2023a), which is computed as
@@
where $\lambda_{1}=0.005$. The final training loss is $\mathcal{L}=\mathcal{L}_{recon}+\mathcal{L}_{SSIM}$.
@@ 4.2. Fast Differentiable Complex-Valued Rasterizer
For the forward pass, we adapt tile-based rendering to track both real and imaginary components of Gaussians. Each thread block collaboratively loads Gaussian data into shared memory, with threads processing individual pixels by accumulating complex field contributions using fast trigonometric operations. Multi-plane rendering is handled through plane-specific filtering directly in the kernel, skipping Gaussians with low assignment probabilities to minimize thread divergence while supporting depth-dependent complex field accumulation. Per pixel, we track final transmittance and the last contributing Gaussian’s position per plane for correct backward gradients.
The backward pass maintains the same tile-based structure but traverses Gaussians back-to-front, reusing the sorted array from the forward pass with pixels processing only up to their recorded last contributor. This ensures constant memory overhead regardless of scene complexity. We store only the final accumulated opacity per pixel during forward pass, then recover intermediate opacity values during the backward pass by dividing the final opacity by each Gaussian’s contribution, eliminating the need for long opacity lists while enabling accurate gradient computation. This implementation efficiently renders complex-valued holographic radiance fields while retaining the performance benefits of the original tile-based rendering structure. For details of the forward operation and backward gradient computation, please refer to Supplementary Sec. C.
4.3. Geometrically Aligned Phase Modeling
Traditional Methods.
A limitation in traditional methods is their dependency on optimizing the complex field for a fixed geometric relationship between the hologram and the 3D scene content. In these methods, a hologram plane $P_{0}(x,y)$ is designed to encode the entire 3D scene as viewed from a specific geometric configuration, with the complex field at the viewing plane described as:

### v2

@@ 3. Background and Problem Definition
Before explaining our approach, we first introduce the key concepts and terminology in this paper to define the research problem we are trying to solve. These concepts are: 3D Gaussians for scene representations, hologram planes for wave recording, and camera parameters for view-dependent rendering.
@@
In line with the approach of Kerbl  et al. (2023), we learn geometry and radiance of scenes through the optimization of 3D Gaussian primitives. We consider a 3D region in Euclidean space, $\Omega\subset\mathbb{R}^{3}$. A scene inside $\Omega$ is described using a set of $N$ oriented 3D Gaussians, and each Gaussian is denoted by $\mathcal{G}_{n}$. Typically, each $\mathcal{G}_{n}$ is parameterized by $k=(\mathbf{c}_{n},\mathbf{x}_{n},\mathbf{R}_{n},\mathbf{S}_{n},\boldsymbol{\alpha}_{n})$, where $\mathbf{c}_{n}\in\mathbb{R}^{3}$ represents color, $\mathbf{x}_{n}\in\mathbb{R}^{3}$ represents the center point of a Gaussian, $\mathbf{R}_{n}\in\mathbb{R}^{4}$ describes the rotation of a Gaussian, $\mathbf{S}_{n}$ represents scale along principal axes defined by $\mathbf{R}_{n}$, and $\boldsymbol{\alpha}_{n}$ represents the opacity. The geometrical shape of a 3D Gaussian is defined as
(2) $\mathcal{G}(\mathbf{x},\mathbf{R},\mathbf{S})\;=\;\exp\Bigl(-\tfrac{1}{2}\,\mathbf{x}^{\top}\,\Sigma^{-1}\,\mathbf{x}\Bigr),$
where the covariance, $\Sigma$, is decomposed into a rotation matrix $\mathbf{R}$, and scaling matrix $\mathbf{S}$, following
(3) $\Sigma\;=\;\mathbf{R}\,\mathbf{S}\,\mathbf{S}^{\top}\,\mathbf{R}^{\top}.$
We project these 3D Gaussians onto 2D planes to render images at a camera plane, a process also known as the differential splatting. The process of differential splatting involves applying the transformation matrix, $\mathbf{W}$, and the Jacobian projection function, $\mathbf{J}$, mapping 3D points to 2D image plane. The covariance in camera-space, $\Sigma^{\prime}$, is defined as
@@
For each pixel, the color and opacity contributed by each Gaussian are computed by evaluating Eq. (2) under its learned parameters. The blending of $N$ ordered Gaussians overlapping a given pixel is defined as follows
@@
We define the hologram plane as $P\in\mathbb{C}^{n_{x}\times n_{y}}$, where $n_{x},n_{y}\in\mathbb{Z}$. The center of $P$ is co-located with a camera pointing at the center of $\Omega$ located at $d$ distance away. We always position $P$ based on the camera pose, location and angles, $r=(x_{c},y_{c},z_{c},\phi_{x},\phi_{y},\phi_{z})$, in Euler space. Each complex pixel $P_{i}$ has a pixel pitch, $\Delta x$. Given input $q=(r,\Delta x,n_{x},n_{y})$, we compute the hologram plane by propagating the complex field rendered from Gaussian primitives to $P(q)$,
(6) P(q)=U(q)h,P(q)=U(q)*h,
where $U(q)$ denotes the complex field rendered from the Gaussian primitives under camera parameters $q$, and $h$ denotes spatial-domain convolutional kernel (Matsushima and Shimobaba, 2009; Zheng et al., 2024). Note that this propagation is an intermediate, differentiable rendering step. Our final goal is to optimize the Gaussian parameters such that the reconstructed intensities after inverse propagation match the focal-stack. A common propagation method is the band-limited ASM, where the transfer function $H_{z}(f_{x},f_{y})$ encodes propagation by a distance $z$ in the frequency domain and its inverse Fourier transform yields the spatial-domain kernel $h$,
@@
In this paper, we also utilize band-limited ASM as the propagation method for complex-valued holographic radiance fields modeling.
@@ 4. Method
In this section, we present our method using the terminology established in Sec. 3. Whereas classical radiance is a scalar energy measure $\mathrm{W/(sr\cdot m^{2})}$ neglecting phase, our approach represents a complex-valued field, capturing both amplitude and phase information to model intensity, interference and diffraction. We use the term complex-valued holographic radiance fields, preserving conceptual links to prior 3D scene representations such as NeRF and 3DGS while reinterpreting radiance within the wave-optics framework. By embedding coherent light modeling into the established radiance field paradigm, our formulation attempts to bridge graphics and holographic rendering.
@@ 4.1. Complex-Valued Holographic Radiance Fields
We provide an overview of our method in Fig. 3. Given a camera pose, $r$, and a target hologram plane, $P$, we partition $\Omega$ into multiple planes in depth. At each depth plane, we render our learnable complex-valued 3D Gaussians into corresponding 2D projections. During projection, we utilize learnable probabilities for our 3D Gaussians that help us determine if a 3D Gaussian contributes to that selected depth plane. Note that learning these probabilities in terms of their contributions at each depth plane helps to inherently capture depth characteristics of a given scene. Once we determine the appropriate 2D projections for each layer, we propagate the complex field from each layer towards our camera to construct our $P$.
@@
For every 3D Gaussian $\mathcal{G}_{n}$ with center $\mathbf{x}_{n}$, we project its 2D projections to plane $\Pi_{l}$ based on learned assignment probabilities. Let $\{\Pi_{1},\Pi_{2},\dots,\Pi_{L}\}$ be a set of discretized parallel planes in the volume, $\Omega$, positioned periodically along the optical axis in front of the view-dependent camera. Each plane $\Pi_{l}$ is separated from its neighbors by a fixed depth interval of $\Delta z$. The planes span a volume of depth and the center of the volume is placed at a propagation distance, $d$, from the camera plane. In this arrangement, the entire depth is symmetrically distributed as in the common literature, ranging from: $Z_{1}=d-\frac{(L-1)}{2}\Delta z$ to $Z_{L}=d+\frac{(L-1)}{2}\Delta z.$ Each plane $\Pi_{l}$ is sampled on a 2D grid that matches the physical resolution of a SLM.
@@
Consider a multi-plane case where Gaussians are projected onto $L$ numbers of planes $\{\Pi_{l}\}_{l=1}^{L}$, we compute each 2D projection of 3D Gaussian following Eq. (4),
(8) $\mathcal{G}^{\text{proj}}(\mathbf{x},\mathbf{R},\mathbf{S},\mathbf{J},\mathbf{W},u)=\exp\left(-\tfrac{1}{2}(u-\mu)^{\top}\Sigma^{\prime-1}(u-\mu)\right),$
where $u=(n_{x},n_{y})$ and $\mu=\mathbf{J}\,\mathbf{W}\,\mathbf{x}$. To model complex-valued holographic radiance fields, we have to modify the original definition of the 3D Gaussian primitive. Specifically, each Gaussian primitive is now parameterized by
(9) $k=(\mathbf{c}_{n},\mathbf{x}_{n},\mathbf{R}_{n},\mathbf{S}_{n},\boldsymbol{\alpha}_{n},\underline{\boldsymbol{\varphi}_{n}},\underline{\boldsymbol{\rho}_{n}}),$
where we add two extra parameters: $\boldsymbol{\varphi}_{n}\in\mathbb{R}^{3}$ to represent inherent phase values across wavelengths, and $\boldsymbol{\rho}_{n}\in\mathbb{R}^{L}$ to represent the plane assignment probabilities, where each element lies in $[0,1]$. Under this formulation, $\mathbf{c}_{n}$ denotes the inherent wave amplitude, replacing its original interpretation as color. The complex field of each 2D projection $U_{n}$ then is given by
@@
The wavenumber term is implicitly included in the phase parameter $\boldsymbol{\varphi}_{n}$, which represents the accumulated phase (including wavelength dependence) at the Gaussian level. When $L=1$, plane assignment reduces to a single plane, the $\boldsymbol{\rho}_{n}$ trivially equals 1.0, and the complex-valued holographic radiance field will generate 2D holograms. For the multi-plane case, we convert the raw, pre-activation logits $\boldsymbol{\rho}^{\prime}_{n}$ to actual probabilities via arg-max followed by one-hot encoding, resulting in a hard assignment vector $\boldsymbol{\rho}_{n}$
(11) $\boldsymbol{\rho}_{n,l}=\mathrm{OneHot}(\arg\max_{l}\,\boldsymbol{\rho}^{\prime}_{n})=\begin{cases}1,&\text{if }l=\displaystyle\arg\max_{l}\,\boldsymbol{\rho}^{\prime}_{n},\\ 0,&\text{otherwise},\end{cases}$
where $\boldsymbol{\rho}_{n,l}$ is the assignment probability of the $n$-th Gaussian to the $l$-th depth plane. However, the argmax operation is non-differentiable, which impedes gradient-based training. To address this, we employ the Straight-Through Estimator (STE) that allows gradient propagation through the discrete assignment
(12) Forward: $\boldsymbol{\rho}_{n}=\text{OneHot}(\underset{l}{\arg\max}\,\boldsymbol{\rho}^{\prime}_{n}),$
Backward: $\frac{\partial\mathcal{L}}{\partial\boldsymbol{\rho}^{\prime}_{n}}=\frac{\partial\mathcal{L}}{\partial\boldsymbol{\rho}_{n}}\cdot\text{softmax}(\boldsymbol{\rho}^{\prime}_{n}/\tau),$
where $\tau$ is the temperature controlling the sharpness of the softmax approximation, in our case, we set $\tau=0.001$, which makes the softmax function close to argmax operation. The STE is widely used in deep learning to approximate gradients through discrete operations (Yin et al., 2019), but it is not a standard component of neural rendering pipelines. We adopt it to our method to circumvent gradient propagation issues.
@@
The final complex field at each plane $\Pi_{l}$ is then defined as
(13) $U_{\Pi_{l}}=\sum_{n=1}^{N}\boldsymbol{\rho}_{n,l}U_{n}\boldsymbol{\alpha}_{n}\prod_{j=1}^{n-1}\bigl(1-\boldsymbol{\alpha}_{j}\bigr)\boldsymbol{\rho}_{j,l}.$
This formulation ensures each complex-valued 3D Gaussian contributes exclusively to its assigned plane. Here, the opacity $\alpha_{n}$ modulates the amplitude of the emitted complex wavefront, representing electromagnetic field strength rather than color blending. It controls the wave emission intensity of each Gaussian and thereby indirectly shapes the resulting interference pattern, while the transmittance product $\prod_{j=1}^{n-1}(1-\alpha_{j})$ accounts for wavefront attenuation due to occlusion by preceding Gaussians. The learned $\boldsymbol{\rho}_{n}$ values guide assignment distribution across view-dependent plane configurations at different viewpoints, minimizing reconstruction loss during training. Unlike the alpha blending used in intensity-based approaches, which indiscriminately mixes RGB colors across all elements, our method creates plane-specific visibility by blending complex numbers (containing both amplitude and phase information) only among Gaussians assigned to the same depth plane to each pixel. Once each layer $\Pi_{l}$ is populated by assigned 2D projections, we use the transfer function $H$ to propagate $\Pi_{l}$ towards the hologram plane $P$, a process we named Forward Recording in our framework. This wave-based propagation extends the traditional intensity neural rendering in Eq. (5). Specifically, for every $\Pi_{l}$, a $(n_{x}^{(l)}\times n_{y}^{(l)})$ grid of complex samples, we compute
@@
where $Z_{l}$ is the distance between depth plane $\Pi_{l}$ and the hologram plane $P$. The forward propagated field $U_{\Pi_{l}\to P}$ is then added with the contributions from the other planes to record the final complex 3D hologram (Shimobaba et al., 2009),
@@
Since each plane is a 2D raster, the computation cost scales linearly with the number of layers $L$ and is dominated by FFTs of size $n_{x}^{(l)}\times n_{y}^{(l)}$. After obtaining $P$, we propagate it back to each depth plane. The back-propagated complex field $U_{P\to\Pi_{l}}$ at depth plane $\Pi_{l}$ is then computed as
@@
The supervision consists of focal-stack intensity images rendered from captured or synthetic multi-view data (Kavaklı et al., 2023a), serving as ground-truth observations and are not produced by any learned baseline. To further improve the image quality of the defocus region, we utilize two complementary loss functions. First, the reconstruction loss $\mathcal{L}_{recon}$ introduced by Kavakli et al. (2023a), computed as
@@
where $\lambda_{1}=0.005$. The final training loss is $\mathcal{L}=\mathcal{L}_{recon}+\mathcal{L}_{SSIM}$. Neither the complex hologram values nor the plane-assignment probabilities are directly supervised in our method, as no GT hologram or plane label per Gaussian is available. Both variables are learned implicitly through reconstruction loss on focal-stack intensities.
View-Dependent Plane Configuration.
At each viewpoint, the depth planes $\{\Pi_{l}\}_{l=1}^{L}$ are always positioned parallel to the hologram plane $P$, ensuring standard band-limited ASM remains valid. To our knowledge, there is not a work adopting off-axis beam propagation in this context, including the prior work (Choi et al., 2025). For novel views with camera pose $r^{\prime}$, we recompute projection matrices $\mathbf{J}^{\prime}$ and $\mathbf{W}^{\prime}$ to project Gaussians onto the reconfigured planes. The plane assignment probabilities $\boldsymbol{\rho}_{n}$ are learned across training views, serving as computational latent variables that guide Gaussian distribution across depth planes.
@@ 4.2. Fast Differentiable Complex-Valued Rasterizer
For the forward pass, we adapt tile-based rendering to track both real and imaginary components of Gaussians. Each thread block collaboratively loads Gaussian data into shared memory, with threads processing individual pixels by accumulating complex-valued field contributions using fast trigonometric operations. Multi-plane rendering is handled through plane-specific filtering directly in the kernel, skipping Gaussians with low assignment probabilities to minimize thread divergence while supporting depth-dependent complex field accumulation. Per pixel, we track final transmittance and the last contributing Gaussian’s position per plane for correct backward gradients.
The backward pass maintains the same tile-based structure but traverses Gaussians back-to-front, reusing the sorted array from the forward pass with pixels processing only up to their recorded last contributor. This ensures constant memory overhead regardless of scene complexity. We store only the final accumulated opacity per pixel during forward pass, then recover intermediate opacity values during the backward pass by dividing the final opacity by each Gaussian’s contribution, eliminating the need for long opacity lists while enabling accurate gradient computation. This implementation efficiently renders complex-valued holographic radiance fields while retaining the performance benefits of the original tile-based rendering structure. For details of the forward operation and backward gradient computation, please refer to Supplementary Sec. D.
4.3. Scene Geometry-Aware Amplitude and Phase Representations
Existing CGH Methods.
A limitation in existing CGH methods is their dependency on optimizing the complex field for a fixed geometric relationship between the scene geometry and the hologram plane. In these methods, a hologram plane $P_{0}(x,y)$ is designed to encode the entire 3D scene when viewed from a specific viewpoint, with the complex field at the viewing plane described as: